# TRACED: Transition-aware Regret Approximation with Co-learnability for Environment Design

### v3

@@ A.4 Ablation Study on the Number of Workers
We analyze the effect of the number of PPO workers, which determines how many agents collect trajectories in parallel. The only difference between our experimental setup and ACCEL (Parker-Holder et al., 2022) is the worker count: we use 16 workers for MiniGrid and 4 for BipedalWalker, whereas the original ACCEL paper uses 32 and 16, respectively. To ensure this configuration does not unfairly disadvantage TRACED, we additionally evaluate MiniGrid with 32 workers. Even under this setting, TRACED consistently outperforms ACCEL, indicating that our gains are robust to the PPO worker configuration (Table 7).
@@ Appendix G Overall Workflow
Algorithm 1 summarizes our overall UED procedure under the TRACED framework. We followed ACCEL (Parker-Holder et al., 2022) in all respects except the procedure used to compute task priority. At each iteration, with probability $1-p_{\mathrm{replay}}$ the teacher enters the exploration phase by sampling a fresh level $\theta\sim\mathcal{G}$, executing the current policy $\pi_{\phi}$ to collect a trajectory, and computing its approximated regret
@@ H.1 Minigrid Environment
Figure 20: MiniGrid zero-shot environments used for evaluation. In SmallCorridor and LargeCorridor, the goal can appear in any of the corridor paths. SimpleCrossing and FourRooms are adapted from MiniGrid (Chevalier-Boisvert et al., 2023), and PerfectMaze from REPAIRED (Jiang et al., 2021a).
The partially observable navigation environment is designed as a $15\times 15$ grid maze, based on MiniGrid (Chevalier-Boisvert et al., 2023). Each tile in the maze can be either an empty tile, a wall, a goal, or the agent itself. The empty tile is a navigable space through which the agent can move, whereas the wall is an impassable obstacle that blocks the agent’s movement. At the beginning of each episode, the initial position of the agent, the goal location, and the wall layout are randomly initialized. Up to 60 walls can be placed throughout the maze, increasing navigational complexity. At each timestep, the agent receives a $5\times 5$ local observation, along with its current direction. The action space is defined as a discrete set of seven possible actions, although only three, turn left, turn right, and move forward, are relevant to maze navigation. These actions are selected without masking the unused action outputs. A reward is provided when the agent successfully reaches the goal, calculated as $1-T/T_{\text{max}}$, where $T$ is the number of steps taken and $T_{\text{max}}=250$ denotes the maximum allowed steps per episode. If the agent fails to reach the goal within the time limit, it does not receive a reward.
To evaluate performance in partially observable navigation, we include additional results on the MiniGrid environments, covering a suite of challenging zero-shot tasks from prior UED work (Dennis et al., 2020; Jiang et al., 2021a). Figure 20 displays all evaluation tasks.
@@ H.2 BipedalWalker Environment
The environment is based on the BipedalWalkerHardcore environment of the OpenAI gym (Brockman et al., 2016). At each timestep, the agent receives a 24-dimensional observation, which includes information such as the hull angle and angular velocity, joint angles and speeds for the hips and knees, ground contact indicators for each foot, the robot’s horizontal and vertical velocities, and lidar-based distance measurements to the ground ahead. The action space is represented by a 4-dimensional vector that controls the torques applied to the robot’s two hip joints and two knee joints. Rewards are structured to encourage efficient and stable locomotion across the terrain. The agent receives positive rewards for processing while maintaining a straight hull posture, with small penalties applied for energy expenditure to discourage unnecessary motor use. The maximum achievable score is $300$. If the agent falls or moves backward, the episode ends immediately, and a penalty of -100 is applied. The maximum episode length is 2000 timesteps. To introduce additional difficulty and test the robustness of the agent’s locomotion policy, the environment includes a variety of challenging terrain features. These include stairs, which consist of sequences of elevated steps that the agent must ascend or descend; pit gaps, which are horizontal gaps over which the agent must jump or step over; stumps, which are vertical obstacles with varying heights and widths that obstruct movement; and surface roughness, which introduces uneven ground textures that can disrupt balance and foot placement.
@@ H.3 Network Structures
For partially observable navigation, the LSTM-based transition prediction model $f_{\phi}$ consists of three components: 1. an image encoder that compresses the input observation $s_{t}$ into a 128‑dimensional latent embedding, 2. an LSTM module (hidden size 128, 1 layer) that processes the concatenation of this embedding and the action $a_{t}$, 3. and an image decoder that reconstructs the predicted next observation $\hat{s}_{t+1}$ from the LSTM output. We adopt an LSTM here because recurrent architectures have proven effective at capturing temporal dependencies in transition and video‑prediction tasks under partial observability (Ha & Schmidhuber, 2018; Oh et al., 2015).
@@
We use L1 loss for MiniGrid because it preserves sharp edges and reduces blurring in image reconstruction (Zhao et al., 2016), while mean squared error is a standard, reliable choice for continuous state regression in locomotion domains (Mnih et al., 2015).
We adopt ACCEL’s network architecture (Parker-Holder et al., 2022), differing only in the LSTM design choice. The PPO network serves as the student agent in our system, directly learning to optimize both the policy and value functions through interaction with the environment. The architecture follows the standard actor-critic structure where the actor head outputs an action distribution, and the critic head predicts the scalar value $V(s)$.
@@ I.2 Difference with CENIE
Unsupervised Environment Design (UED) defines a co-evolutionary framework in which a teacher agent generates task instances for a student policy, selecting those tasks with high learning potential to promote robust generalization to unseen environments (Wang et al., 2019; 2020; Dennis et al., 2020; Chung et al., 2024). PAIRED (Dennis et al., 2020) trains the teacher to maximize the regret between a protagonist and antagonist policy, causing the teacher to create increasingly challenging yet solvable environments. However, non-stationary task distributions and the high dimensionality of the task space impede convergence to optimal curricula. To mitigate these limitations, Mediratta et al. (2023) introduce entropy bonuses, alternative optimization algorithms, and online behavioral cloning, and Clutr (Azad et al., 2023a) employs variational autoencoder-based unsupervised task representation learning to compress the task manifold and yield more stable curricula.
@@
Another line of work is based on Prioritized Level Replay (PLR) (Jiang et al., 2021b), which avoids training a teacher agent and instead randomly generates tasks while introducing a replay buffer that stores previously generated tasks. This approach prioritizes tasks with high learning potential from the buffer to more effectively train the student agent. PLR  (Jiang et al., 2021a) make improvement of PLR, which used stop-gradient on trajectories from newly generated tasks but updated only with replayed tasks from buffer. This leads to better performance on unseen environments counterintuitively. ACCEL (Parker-Holder et al., 2022) extends the PLR by implementing an evolutionary approach that makes small mutation to high-regret levels from the buffer. This evolutionary mechanism enables the development of progressively more complex challenges starting with simpler tasks. However, these methods (Jiang et al., 2021b; a; Parker-Holder et al., 2022) approximate regret solely via value loss, thereby implicitly learn true environment dynamics that could improve regret estimation.
@@
Moving beyond regret-based approaches, CENIE (Teoh et al., 2024) integrates novelty with regret to expose agents to more diverse learning situations. Using Gaussian Mixture Models, CENIE quantifies how much a new task’s trajectory differs from previous experiences stored in the buffer. This curriculum-awareness leads to better exploration and generalization. DIPLR (Li et al., 2023) introduces a diversity that quantifies the similarity between different levels by computing the Wasserstein distance between their occupancy distributions over state-action trajectories. MBeDED (Li et al., 2024) introduces the marginal benefit metric to quantify the performance gain of a student policy from training on a generated task. It compares a base policy (before training) and the updated student policy by measuring the difference in their expected returns on that task. So far, to the best of our knowledge, there are no prior works that explicitly consider Co-Learnability between tasks.
@@
UED’s benefit on robust generalization performance in unseen environments has led to its application in other domains. For Multi-Agent Reinforcement Learning, MAESTRO (Samvelyan et al., 2023) extends UED to competitive multi-agent settings by creating a regret-based curriculum over the joint environment and co-player space, leveraging self-play to generate opponents while training agents robust to environment and co-player variations. In contrast, RACCOON (Erlebach & Cook, 2024) applies UED to cooperative scenarios, developing a curriculum that prioritizes high-regret partners from pre-trained partners to enhance collaboration. For meta-RL, GROOVE (Jackson et al., 2023) suggested policy meta-optimization (PMO) introducing algorithmic regret (AR), which measures the performance gap between a meta-learned optimizer and RL algorithms like A2C, creating curricula that identify informative environments for meta-training. DIVA (Costales & Nikolaidis, 2024) is the first method that extends UED to semi-supervised environment design, employing quality-diversity search to maintain a population of diverse training tasks in open-ended simulators, maximizing behavioral coverage. In sim-to-real settings, GCL (Wang et al., 2024) aligns the simulated curriculum with actual deployment tasks by sampling from real-world task distributions and adapting subsequent task generation based on the robot’s past performance.
@@
Similar to UED works, Curriculum Reinforcement Learning (CRL) accelerates agent training by constructing a sequence of progressively challenging tasks rather than directly confronting the agent with a single, complex target task (Narvekar et al., 2020). Early approaches distinguished between fixed curricula in which task order is predefined and self-paced learning, in which task selection adapts dynamically based on the agent’s learning progress (Klink et al., 2020). Modern CRL methods have advanced this paradigm by framing curriculum generation as an inference process (Eimer et al., 2021; Klink et al., 2022). WAKER (Rigter et al., 2024) leverages reward-free curricula to train robust world models, thereby underscoring the versatility of self-paced approaches in diverse learning scenarios. However, these approaches assume access to the target task distribution, which conflicts with UED’s core assumption of an unknown distribution.
@@
Robust PLR (PLR(Jiang et al., 2021a) was introduced as an extension of the original PLR (Jiang et al., 2021b). While DR treats all tasks equally, PLR prioritizes tasks with high learning potential, maintaining a buffer $\Lambda$ of previously encountered tasks for replay. When sampling new tasks, the agent’s policy parameters are stop-gradiented, meaning the policy is not updated on trajectories collected from these newly sampled tasks. The collected trajectories are used to compute PVL scores to decide whether to add the task to the replay buffer. Policy updates are exclusively performed on tasks sampled from the replay buffer. By stopping the gradient on randomly sampled levels, PLR ensures that the policy is only updated on tasks specifically selected to maximize regret, leading to more robust generalization compared to both DR and PLR unintuitively.
@@
Adversarially Compounding Complexity by Editing Levels (ACCEL) (Parker-Holder et al., 2022) actively evolves environments through an editing mechanism, allowing it to more efficiently explore the environment design space. This enables reusing the structure of sampled levels in the buffer for high-regret, rather than curating randomly sampled levels for high-regret. ACCEL maintains a buffer $\Lambda$ and employs a cycle of sampling, editing, and curation. The key insight of ACCEL is that regret serves as a domain-agnostic fitness function for evolution, enabling it to produce batches of levels at the frontier of agent capabilities. The editing mechanism involves making small mutations to previously high-regret levels, which can operate directly on environment elements such as blocks in MiniGrid. This evolutionary process creates an expanding frontier that matches the agent’s capabilities, starting with simple levels and progressively increasing in complexity. Like PLR, ACCEL employs stop-gradient when evaluating new or edited levels to ensure theoretical guarantees.
@@
Coverage-based Evaluation of Novelty In Environment (CENIE) (Teoh et al., 2024) introduces environment novelty as a complementary objective to regret in UED. CENIE quantifies environment novelty through state-action space coverage derived from the agent’s accumulated experiences across previous environments in its curriculum. The framework operates on the intuition that a novel environment should induce unfamiliar experiences, pushing the student agent into unexplored regions of the state-action space. At the core of CENIE’s implementation is the use of Gaussian Mixture Models (GMMs) to model the distribution of state-action pairs from the agent’s past experiences. Given a state-action buffer $\Gamma$ containing pairs collected from previous environments, CENIE fits a GMM with parameters $\lambda_{\Gamma}$ to represent this distribution. The novelty of a candidate environment is then quantified by measuring the similarity between its newly observed state-action pairs and the learned distribution of past state-action experiences. CENIE integrates with existing UED algorithms by combining both novelty and regret into a unified priority score for environment selection.
@@
Adversarial Environment Design via Regret-Guided Diffusion Models (ADD) (Chung et al., 2024) frames environment generation as a diffusion process steered by the agent’s regret. Starting from uniformly sampled tasks, a pretrained score network $s_{\phi}$ is guided by the gradient of a differentiable regret estimate, computed via a learned return critic, during reverse diffusion.
@@
In our empirical study, DR, PLR, and ACCEL were evaluated with the DCD codebase (Jiang et al., 2022), and assessed ADD with the ADD codebase (Chung et al., 2024). CENIE’s performance data were incorporated from its original publication, as no open-source implementation was available.
@@
Extension to other RL algorithms. Our experiments have been conducted with PPO-based student agents. While this allows for fair comparisons with previous work, there is a much broader range of algorithms that may interact differently with our regret-based curriculum. Off-policy methods such as SAC (Haarnoja et al., 2018) and TD3 (Fujimoto et al., 2018) feature distinct exploration strategies, which may yield different performance dynamics under our curriculum framework. Moreover, Model-based approaches like DreamerV3 (Hafner et al., 2025) build world models for planning, offering complementary insights into curriculum design. Evaluating these methods under our curriculum framework remains an important direction for future work.
Extend to other UED methods. We empirically demonstrated that TRACED yields more effective curricula, thereby enhancing agents’ generalization capabilities at half the training cost. While our experiments focus on the ACCEL framework, future work will integrate our approach into alternative UED methods (PLR (Jiang et al., 2021a) and CENIE (Teoh et al., 2024)) and conduct comprehensive experiments to assess its performance in these settings.
Evaluation in other domains. Our experimental validation has been conducted in MiniGrid (Chevalier-Boisvert et al., 2023; Dennis et al., 2020) and BipedalWalker (Brockman et al., 2016; Parker-Holder et al., 2022). Evaluating TRACED in other domains such as Car Racing (Brockman et al., 2016; Jiang et al., 2021a), physical robotics applications (Mahmood et al., 2018), and reasoning task ARC (Chollet, 2019) would provide a more comprehensive assessment of the approach’s generality and scalability across diverse settings.
@@
Table 10 summarizes the hyperparameters used. Unless otherwise noted, values are adopted directly from the DCD repository (Jiang et al., 2022). The only deviation from the DCD defaults is the number of PPO workers (see Appendix A.4). We tune two key weights, $\alpha$ and $\beta$, on MiniGrid and BipedalWalker, exploring $\alpha\in\{0.5,1.0,1.5\}$ and $\beta\in\{0.6,0.8,1.0\}$.
@@
Table 11 reports per-task zero-shot solved rates on twelve held-out MiniGrid mazes, averaged over 10 seeds (mean $\pm$ s.e.), following Agarwal et al. (2021). We compare TRACED at 10k updates against DR, PLR, ACCEL, and ADD (each at 10k and 20k). At 10k, TRACED achieves the highest solved rate in six mazes (16Rooms2, SimpleCrossing, Labyrinth, Labyrinth2, Maze, PerfectMaze(M)) and the second-best in one maze (Maze2). Averaged across all twelve mazes, TRACED attains a mean solved rate of 0.66, outperforming all baseline configurations.
@@
Table 14 compares TRACED’s zero-shot returns with DR, PLR, ADD, ACCEL, and ACCEL-CENIE on six BipedalWalker terrains, averaged over five seeds (mean $\pm$ s.e.). ACCEL-CENIE is estimated by applying the ACCEL-to-CENIE performance ratio reported in Teoh et al. (2024) to our ACCEL implementation. At 10k updates, TRACED achieves the highest mean return on three terrains (Basic, Hardcore, and Stump) and attains an overall mean of $89.95\pm 31.72$, outperforming all baselines.

### v4

@@ A.4 Ablation Study on the Number of Workers
We analyze the effect of the number of PPO workers, which determines how many agents collect trajectories in parallel. The only difference between our experimental setup and ACCEL (ACCEL) is the worker count: we use 16 workers for MiniGrid and 4 for BipedalWalker, whereas the original ACCEL paper uses 32 and 16, respectively. To ensure this configuration does not unfairly disadvantage TRACED, we additionally evaluate MiniGrid with 32 workers. Even under this setting, TRACED consistently outperforms ACCEL, indicating that our gains are robust to the PPO worker configuration (Table 7).
@@ Appendix G Overall Workflow
Algorithm 1 summarizes our overall UED procedure under the TRACED framework. We followed ACCEL (ACCEL) in all respects except the procedure used to compute task priority. At each iteration, with probability $1-p_{\mathrm{replay}}$ the teacher enters the exploration phase by sampling a fresh level $\theta\sim\mathcal{G}$, executing the current policy $\pi_{\phi}$ to collect a trajectory, and computing its approximated regret
@@ H.1 Minigrid Environment
Figure 20: MiniGrid zero-shot environments used for evaluation. In SmallCorridor and LargeCorridor, the goal can appear in any of the corridor paths. SimpleCrossing and FourRooms are adapted from MiniGrid (MinigridMiniworld23), and PerfectMaze from REPAIRED (REPAIRED).
The partially observable navigation environment is designed as a $15\times 15$ grid maze, based on MiniGrid (MinigridMiniworld23). Each tile in the maze can be either an empty tile, a wall, a goal, or the agent itself. The empty tile is a navigable space through which the agent can move, whereas the wall is an impassable obstacle that blocks the agent’s movement. At the beginning of each episode, the initial position of the agent, the goal location, and the wall layout are randomly initialized. Up to 60 walls can be placed throughout the maze, increasing navigational complexity. At each timestep, the agent receives a $5\times 5$ local observation, along with its current direction. The action space is defined as a discrete set of seven possible actions, although only three, turn left, turn right, and move forward, are relevant to maze navigation. These actions are selected without masking the unused action outputs. A reward is provided when the agent successfully reaches the goal, calculated as $1-T/T_{\text{max}}$, where $T$ is the number of steps taken and $T_{\text{max}}=250$ denotes the maximum allowed steps per episode. If the agent fails to reach the goal within the time limit, it does not receive a reward.
To evaluate performance in partially observable navigation, we include additional results on the MiniGrid environments, covering a suite of challenging zero-shot tasks from prior UED work (PAIRED; REPAIRED). Figure 20 displays all evaluation tasks.
@@ H.2 BipedalWalker Environment
The environment is based on the BipedalWalkerHardcore environment of the OpenAI gym (Brockman2016Gym). At each timestep, the agent receives a 24-dimensional observation, which includes information such as the hull angle and angular velocity, joint angles and speeds for the hips and knees, ground contact indicators for each foot, the robot’s horizontal and vertical velocities, and lidar-based distance measurements to the ground ahead. The action space is represented by a 4-dimensional vector that controls the torques applied to the robot’s two hip joints and two knee joints. Rewards are structured to encourage efficient and stable locomotion across the terrain. The agent receives positive rewards for processing while maintaining a straight hull posture, with small penalties applied for energy expenditure to discourage unnecessary motor use. The maximum achievable score is $300$. If the agent falls or moves backward, the episode ends immediately, and a penalty of -100 is applied. The maximum episode length is 2000 timesteps. To introduce additional difficulty and test the robustness of the agent’s locomotion policy, the environment includes a variety of challenging terrain features. These include stairs, which consist of sequences of elevated steps that the agent must ascend or descend; pit gaps, which are horizontal gaps over which the agent must jump or step over; stumps, which are vertical obstacles with varying heights and widths that obstruct movement; and surface roughness, which introduces uneven ground textures that can disrupt balance and foot placement.
@@ H.3 Network Structures
For partially observable navigation, the LSTM-based transition prediction model $f_{\phi}$ consists of three components: 1. an image encoder that compresses the input observation $s_{t}$ into a 128‑dimensional latent embedding, 2. an LSTM module (hidden size 128, 1 layer) that processes the concatenation of this embedding and the action $a_{t}$, 3. and an image decoder that reconstructs the predicted next observation $\hat{s}_{t+1}$ from the LSTM output. We adopt an LSTM here because recurrent architectures have proven effective at capturing temporal dependencies in transition and video‑prediction tasks under partial observability (worldmodels; ohjunhyuk2015action).
@@
We use L1 loss for MiniGrid because it preserves sharp edges and reduces blurring in image reconstruction (zhao2016loss), while mean squared error is a standard, reliable choice for continuous state regression in locomotion domains (mnih2015human).
We adopt ACCEL’s network architecture (ACCEL), differing only in the LSTM design choice. The PPO network serves as the student agent in our system, directly learning to optimize both the policy and value functions through interaction with the environment. The architecture follows the standard actor-critic structure where the actor head outputs an action distribution, and the critic head predicts the scalar value $V(s)$.
@@ I.2 Difference with CENIE
Unsupervised Environment Design (UED) defines a co-evolutionary framework in which a teacher agent generates task instances for a student policy, selecting those tasks with high learning potential to promote robust generalization to unseen environments (POET; EnhancedPOET; PAIRED; Chung2024AdversarialDiffusion). PAIRED (PAIRED) trains the teacher to maximize the regret between a protagonist and antagonist policy, causing the teacher to create increasingly challenging yet solvable environments. However, non-stationary task distributions and the high dimensionality of the task space impede convergence to optimal curricula. To mitigate these limitations, Stabilizing introduce entropy bonuses, alternative optimization algorithms, and online behavioral cloning, and Clutr (CLUTR) employs variational autoencoder-based unsupervised task representation learning to compress the task manifold and yield more stable curricula.
@@
Another line of work is based on Prioritized Level Replay (PLR) (PLR), which avoids training a teacher agent and instead randomly generates tasks while introducing a replay buffer that stores previously generated tasks. This approach prioritizes tasks with high learning potential from the buffer to more effectively train the student agent. PLR  (REPAIRED) make improvement of PLR, which used stop-gradient on trajectories from newly generated tasks but updated only with replayed tasks from buffer. This leads to better performance on unseen environments counterintuitively. ACCEL (ACCEL) extends the PLR by implementing an evolutionary approach that makes small mutation to high-regret levels from the buffer. This evolutionary mechanism enables the development of progressively more complex challenges starting with simpler tasks. However, these methods (PLR; REPAIRED; ACCEL) approximate regret solely via value loss, thereby implicitly learn true environment dynamics that could improve regret estimation.
@@
Moving beyond regret-based approaches, CENIE (Jayden2024CENIE) integrates novelty with regret to expose agents to more diverse learning situations. Using Gaussian Mixture Models, CENIE quantifies how much a new task’s trajectory differs from previous experiences stored in the buffer. This curriculum-awareness leads to better exploration and generalization. DIPLR (DIPLR) introduces a diversity that quantifies the similarity between different levels by computing the Wasserstein distance between their occupancy distributions over state-action trajectories. MBeDED (Marginal) introduces the marginal benefit metric to quantify the performance gain of a student policy from training on a generated task. It compares a base policy (before training) and the updated student policy by measuring the difference in their expected returns on that task. SFL (rutherford2024no) assigns a learnability score to each level, computed from the empirical success rate achieved by evaluating the agent on sampled levels. Levels whose success rates fall in the intermediate range, neither too easy nor too difficult, are prioritized, guiding the agent toward tasks that are most conducive to further learning. A related idea appears in GoalGAN (Florensa2018GoalGAN), which is not a UED method but similarly adapts task difficulty by training a generative model to propose goals of suitable challenge. So far, to the best of our knowledge, there are no prior works that explicitly consider Co-Learnability between tasks.
@@
UED’s benefit on robust generalization performance in unseen environments has led to its application in other domains. For Multi-Agent Reinforcement Learning, MAESTRO (samvelyan2023maestro) extends UED to competitive multi-agent settings by creating a regret-based curriculum over the joint environment and co-player space, leveraging self-play to generate opponents while training agents robust to environment and co-player variations. In contrast, RACCOON (erlebach2024raccoon) applies UED to cooperative scenarios, developing a curriculum that prioritizes high-regret partners from pre-trained partners to enhance collaboration. For meta-RL, GROOVE (groove) suggested policy meta-optimization (PMO) introducing algorithmic regret (AR), which measures the performance gap between a meta-learned optimizer and RL algorithms like A2C, creating curricula that identify informative environments for meta-training. DIVA (DIVA) is the first method that extends UED to semi-supervised environment design, employing quality-diversity search to maintain a population of diverse training tasks in open-ended simulators, maximizing behavioral coverage. In sim-to-real settings, GCL (wang2024grounded) aligns the simulated curriculum with actual deployment tasks by sampling from real-world task distributions and adapting subsequent task generation based on the robot’s past performance.
@@
Similar to UED works, Curriculum Reinforcement Learning (CRL) accelerates agent training by constructing a sequence of progressively challenging tasks rather than directly confronting the agent with a single, complex target task (CurriculmRLSurvey). Early approaches distinguished between fixed curricula in which task order is predefined and self-paced learning, in which task selection adapts dynamically based on the agent’s learning progress (SPDRL). Modern CRL methods have advanced this paradigm by framing curriculum generation as an inference process (SPaCE; currot). WAKER (rigter2024rewardfree) leverages reward-free curricula to train robust world models, thereby underscoring the versatility of self-paced approaches in diverse learning scenarios. Akin to the concept underlying SFL (rutherford2024no), ProCuRL (ProximalCurriculum2023) operationalizes the pedagogical notion of the Zone of Proximal Development by using probability of success score-based scoring combined with softmax sampling to prioritize tasks of intermediate difficulty. However, these approaches assume access to the target task distribution, which conflicts with UED’s core assumption of an unknown distribution.
@@
Robust PLR (PLR(REPAIRED) was introduced as an extension of the original PLR (PLR). While DR treats all tasks equally, PLR prioritizes tasks with high learning potential, maintaining a buffer $\Lambda$ of previously encountered tasks for replay. When sampling new tasks, the agent’s policy parameters are stop-gradiented, meaning the policy is not updated on trajectories collected from these newly sampled tasks. The collected trajectories are used to compute PVL scores to decide whether to add the task to the replay buffer. Policy updates are exclusively performed on tasks sampled from the replay buffer. By stopping the gradient on randomly sampled levels, PLR ensures that the policy is only updated on tasks specifically selected to maximize regret, leading to more robust generalization compared to both DR and PLR unintuitively.
@@
Adversarially Compounding Complexity by Editing Levels (ACCEL) (ACCEL) actively evolves environments through an editing mechanism, allowing it to more efficiently explore the environment design space. This enables reusing the structure of sampled levels in the buffer for high-regret, rather than curating randomly sampled levels for high-regret. ACCEL maintains a buffer $\Lambda$ and employs a cycle of sampling, editing, and curation. The key insight of ACCEL is that regret serves as a domain-agnostic fitness function for evolution, enabling it to produce batches of levels at the frontier of agent capabilities. The editing mechanism involves making small mutations to previously high-regret levels, which can operate directly on environment elements such as blocks in MiniGrid. This evolutionary process creates an expanding frontier that matches the agent’s capabilities, starting with simple levels and progressively increasing in complexity. Like PLR, ACCEL employs stop-gradient when evaluating new or edited levels to ensure theoretical guarantees.
@@
Coverage-based Evaluation of Novelty In Environment (CENIE) (Jayden2024CENIE) introduces environment novelty as a complementary objective to regret in UED. CENIE quantifies environment novelty through state-action space coverage derived from the agent’s accumulated experiences across previous environments in its curriculum. The framework operates on the intuition that a novel environment should induce unfamiliar experiences, pushing the student agent into unexplored regions of the state-action space. At the core of CENIE’s implementation is the use of Gaussian Mixture Models (GMMs) to model the distribution of state-action pairs from the agent’s past experiences. Given a state-action buffer $\Gamma$ containing pairs collected from previous environments, CENIE fits a GMM with parameters $\lambda_{\Gamma}$ to represent this distribution. The novelty of a candidate environment is then quantified by measuring the similarity between its newly observed state-action pairs and the learned distribution of past state-action experiences. CENIE integrates with existing UED algorithms by combining both novelty and regret into a unified priority score for environment selection.
@@
Adversarial Environment Design via Regret-Guided Diffusion Models (ADD) (Chung2024AdversarialDiffusion) frames environment generation as a diffusion process steered by the agent’s regret. Starting from uniformly sampled tasks, a pretrained score network $s_{\phi}$ is guided by the gradient of a differentiable regret estimate, computed via a learned return critic, during reverse diffusion.
@@
In our empirical study, DR, PLR, and ACCEL were evaluated with the DCD codebase (dcd_github), and assessed ADD with the ADD codebase (Chung2024AdversarialDiffusion). CENIE’s performance data were incorporated from its original publication, as no open-source implementation was available.
@@
Extension to other RL algorithms. Our experiments have been conducted with PPO-based student agents. While this allows for fair comparisons with previous work, there is a much broader range of algorithms that may interact differently with our regret-based curriculum. Off-policy methods such as SAC (haarnoja2018soft) and TD3 (fujimoto2018addressing) feature distinct exploration strategies, which may yield different performance dynamics under our curriculum framework. Moreover, Model-based approaches like DreamerV3 (hafner2025) build world models for planning, offering complementary insights into curriculum design. Evaluating these methods under our curriculum framework remains an important direction for future work.
Extend to other UED methods. We empirically demonstrated that TRACED yields more effective curricula, thereby enhancing agents’ generalization capabilities at half the training cost. While our experiments focus on the ACCEL framework, future work will integrate our approach into alternative UED methods (PLR (REPAIRED) and CENIE (Jayden2024CENIE)) and conduct comprehensive experiments to assess its performance in these settings.
Evaluation in other domains. Our experimental validation has been conducted in MiniGrid (MinigridMiniworld23; PAIRED) and BipedalWalker (Brockman2016Gym; ACCEL). Evaluating TRACED in other domains such as Car Racing (Brockman2016Gym; REPAIRED), physical robotics applications (mahmood2018benchmarking), and reasoning task ARC (chollet2019measure) would provide a more comprehensive assessment of the approach’s generality and scalability across diverse settings.
@@
Table 10 summarizes the hyperparameters used. Unless otherwise noted, values are adopted directly from the DCD repository (dcd_github). The only deviation from the DCD defaults is the number of PPO workers (see Appendix A.4). We tune two key weights, $\alpha$ and $\beta$, on MiniGrid and BipedalWalker, exploring $\alpha\in\{0.5,1.0,1.5\}$ and $\beta\in\{0.6,0.8,1.0\}$.
@@
Table 11 reports per-task zero-shot solved rates on twelve held-out MiniGrid mazes, averaged over 10 seeds (mean $\pm$ s.e.), following Agarwal2021Reliable. We compare TRACED at 10k updates against DR, PLR, ACCEL, and ADD (each at 10k and 20k). At 10k, TRACED achieves the highest solved rate in six mazes (16Rooms2, SimpleCrossing, Labyrinth, Labyrinth2, Maze, PerfectMaze(M)) and the second-best in one maze (Maze2). Averaged across all twelve mazes, TRACED attains a mean solved rate of 0.66, outperforming all baseline configurations.
@@
(a) PerfectMazeXL
(b) PerfectMazeLarge
Figure 22: Aggregate zero-shot performance on PerfectMaze. Full metric curves (Median, IQM, Mean, Optimality Gap) for PerfectMazeXL (top) and PerfectMazeLarge (bottom).
Figure 22 presents the full aggregate metrics for both PerfectMaze benchmarks. Despite the high variance inherent to these large-scale tasks, TRACED achieves the strongest 10k performance and closely matches the best 20k performance among all baselines.
@@
Table 14 compares TRACED’s zero-shot returns with DR, PLR, ADD, ACCEL, and ACCEL-CENIE on six BipedalWalker terrains, averaged over five seeds (mean $\pm$ s.e.). ACCEL-CENIE is estimated by applying the ACCEL-to-CENIE performance ratio reported in Jayden2024CENIE to our ACCEL implementation. At 10k updates, TRACED achieves the highest mean return on three terrains (Basic, Hardcore, and Stump) and attains an overall mean of $89.95\pm 31.72$, outperforming all baselines.